OpenStudy (anonymous):

Solve the following using the quadratic formula. Round to the nearest tenth, if necessary. x2 + 5x - 2 = 0

OpenStudy (anonymous):

x is .372281 or -5.37228

OpenStudy (anonymous):

just plug it into the quadratic formula

OpenStudy (anonymous):

x= -b plus or minus the square root of b^2-4ac divided by 2a

OpenStudy (anonymous):

x2 + 5x - 2 = 0 is 1x2 + 5x - 2 = 0 ^ ^ ^ a b c

OpenStudy (anonymous):

x=-5+(25-8)/2 or -5-(25-8)/2 x=-5+17 or -5-17 x=12 or -22 since -22 is negative answer is 12